# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

### v1

@@ 2.1 Problem Statement
$D^{k}=\left\lceil\frac{S^{k}}{T^{k}}\right\rceil-1$. The degeneracy is the start-to-end latency normalized in the number of periods minus 1. Thus, a chain with start-to-end latency shorter than its period has degeneracy zero.
$\alpha$-degeneracy of a chain $k$:
@@
Given an $\alpha\in(0,1]$ $D^{k}(\alpha)=\left\lceil\frac{S^{k}}{\alpha\cdot T^{k}}\right\rceil-1$. This criterion uses finer granularity of the degeneracy whose minimization might be more restrictive: for example, $D^{k}(0.5)=0$, which means that the start-to-end latency of chain $\mathcal{C}^{k}$ fits in half of the chain period.
Notice that the degeneracy of a chain can be computed by considering the core schedule of the chain and the number of postponing operations necessary to satisfy the precedence constraints. So, the degeneracy of the chain is the value of $q_{n^{k}}$ mentioned earlier.
@@
For each measure, we consider the max value and the sum of values:
  • 1.
    max degeneracy $D_{\max}=\max_{k=1,\ldots,\kappa}D^{k}$ and degeneracy $D_{sum}=\sum_{k=1}^{\kappa}D^{k}$.
  • 2.
    similarly max and sum $\alpha$-degeneracy can be defined $D_{\max}({\alpha}),D_{sum}({\alpha})$.
2.2 Special Cases
@@
In the paper, we consider both general $PD|T_{i}^{harm},jit=0,chains|criteria$ instances, as well as instances generated with a set of realistic assumptions (e.g., using underlying network topology) to fit the practical applications where chains represent messages sent over communication links. Some of these assumptions, which will be used in the following sections of the paper, are:
Limited number of resources
@@
Harmonic processing times.
In this case, all values of processing times and periods are together in a harmonic set. This makes the single resource feasibility problem polynomial [6]. Graham notation for such case is $PD|(T_{i},p_{i})^{harm},jit=0,chains|criteria$.
Topology
In this case, the communications links are organized along a topology (line, tree, for example). A chain is characterized by the source resource and the sink resource, and the chain’s tasks are assigned to the sequence of resources on the unique path from the source to the sink.
Equal processing time on a chain
In this case, inspired by the most common case in communications, point-to-point communications have the same duration since the packets have the same length and communication links have the same speed. In such case $p_{i}^{k}=p^{k}$.
Scattering
A special case where all the first tasks of the chains are assigned to the same resource; we call this scattering (as in the context of parallel computing [34]). Proofs related to this case are described in Section 3. Figure 1 shows a scattering example over a line topology with equal processing times.
@@ 2.3 Structural Properties
In [12], we established a formal relation between periodic schedules on a single resource and a particular case of 2D packing problem. The packing view is a crucial inspiration behind the warm-start approach described later.
@@ 2.3.1 Single Resource Core Schedule Coordinates
Consider a single resource $m$ and a feasible core schedule $\sigma$ of the independent periodic tasks of $\mathcal{J}_{m}$. Let $o_{m}$ be the start time of the first task $J_{f}^{k}$ on resource $m$ with the least period $w$. We consider $o_{m}$ as the origin for resource $m$. So we define the shifted core schedule $s_{i}^{k}$ for each task $J_{i}^{k}\in\mathcal{J}_{m}$ such that:
@@
We call time interval $[o_{m},o_{m}+\hat{H})$ the ”observation interval.” Notice that for any integer $v\geq 1$, the schedule $\sigma$ of time interval $[o_{m}+(v-1)\hat{H},o_{m}+v\hat{H})$ on resource $m$ is the same as the schedule of the observation interval shifted by $(v-1)\hat{H}$ time units. This offset may differ for different resources.
@@ 2.3.2 Associated HD2D Packing Problem
To each task $J_{i}^{k}\in\mathcal{J}_{m}$ we associate a rectangle $R_{i}^{k}$ of width $\ell_{i}^{k}=p_{i}^{k}$ and height $0pt{i}{k}=\frac{\hat{H}}{T^{k}}$, to fit in a bin of width $w$ and height $H=\frac{\hat{H}}{w}$ without overlapping rectangles. Moreover, the packing has to be height divisible, which means that the vertical coordinate of each rectangle is a multiple of its height.
(a) Schedule of a single observation interval of resource 2 arranged in rows of length 40
(b) Height-divisible 2D packing of resource 2
Figure 2: Observation interval and packing perspective of a schedule of resource 2 from Fig. 1.
Paper [12] gives the formal correspondence between the start time of a task in the core schedule $s_{i}^{k}$ and coordinates of the bottom left corner of the associated rectangle. To give an intuition of this correspondence, we develop it on an example in Fig. 2. The main idea is to arrange the periodic schedule in rows. A row indexed by $v$ represents the schedule of interval $[vw,(v+1)w)$ of the observation interval.
In Fig. 2(a), the single ”observation interval” associated with the schedule of resource 2 from motivational example Fig. 1 is shown. Ten chains $\mathcal{C}^{1},\dots,\mathcal{C}^{10}$ pass through the resource. $\mathcal{C}^{1},\mathcal{C}^{2}$ have the shortest period $\tau(1)=40$. Chains $\mathcal{C}^{3}-\mathcal{C}^{7}$ have a period equal to 80; the rest have a period equal to 240, which is the hyperperiod. Origin 00 is set by the start time of chain $\mathcal{C}^{1}$, and we stack each interval of length $w=40$ on top of the previous one. So row $v$ describes the schedule of interval $[v\cdot w,(v+1)w)$. Horizontal coordinate $u$ corresponds to the relative position of the task’s occurrence in the interval.
@@
Figure 2(b) shows the corresponding height-divisible packing view of the same solution. Coordinates of Rectangles are described with coordinates of their bottom-left corners $(x,y)$. The packing is obtained by interchanging the rows as follows: we swap the two-digit row numbers in Figure 2(a) such that the most significant digit becomes the least significant digit and vice versa. Finally, we obtain Figure 2(b) by ordering the rows according to their new two-digit numbers. For example, the row with $\mathcal{C}^{6}$ has number $10_{3}1_{2}=$ 1 in Figure 2(a) and after the swap, the same row has number $31_{2}0_{3}=$ 3 in Figure 2(b). We can see the occurrences of the original tasks have been merged, and the dimensions of the resulting rectangles correspond to the rectangles of the associated packing instance. For example, tasks 1-2 with the shortest period are transformed into rectangles spanning the entire height of the bin. Tasks 3-6 merge three of their occurrences into a rectangle with height 3, and finally, rectangles associated with the longest-period tasks 8-10 have a height equal to 1.
@@ 3 Complexity Analysis
In this section, we consider related complexity results to explain the complexity of our problems. First, we know that scheduling periodic tasks with harmonic periods on a single resource is a strongly NP-hard problem [5]. So, minimizing one of our criteria seems at least as difficult as this problem and is still NP-hard unless very strong assumptions are made. In the rest of the section, we show some very special polynomial cases of our problem and present a strongly NP-hard subproblem.
3.1 Scattering on a Line Topology with Two Resources
Let us consider a special case of the line topology with two resources and a scattering hypothesis. So, all chains have their first task on resource 1 and second on resource 2.
3.1.1 Single Period

### v2

@@ 2.2.3 Objectives
The degeneracy is the start-to-end latency normalized in the number of periods minus 1. Thus, a chain with start-to-end latency shorter than its period has degeneracy zero.
Definition 10 ($\alpha$-degeneracy of a chain $k$).
Given $\alpha\in(0,1]$:
$D^{k}(\alpha)\coloneqq\left\lceil\frac{S^{k}}{\alpha\cdot T^{k}}\right\rceil-1.$ (8)
@@
This function uses finer granularity of the degeneracy whose minimization might be more restrictive: for example, $D^{k}(0.5)=0$, which means that the start-to-end latency of chain $\mathcal{C}^{k}$ fits in half of the chain period.
@@
For each measure, we consider the maximum value and the sum of values as objective functions to be minimized:
Definition 11 (Objective functions).
We call maximum degeneracy the objective
$D_{\max}\coloneqq\max_{k=1,\ldots,\mu}D^{k}$ (9)
We call degeneracy the sum of degeneracies of all chains:
$D_{sum}\coloneqq\sum_{k=1}^{\mu}D^{k}$ (10)
Similarly can be defined the maximum and sum $\alpha$ degeneracy $D_{\max}({\alpha}),D_{sum}({\alpha})$.
According to the standard Graham notation extended to handle PSPs in [5], our problem is denoted by ${PD|T_{i}^{harm},jit=0,chains|objective}$. $PD$ stands for parallel dedicated processors, $chains$ for chains of tasks, $T_{i}^{harm}$ for harmonic periods, $jit=0$ since no jitter is allowed (i.e., the starting time of a task is entirely determined by the starting time of its first occurrence and by its period). We recall that all tasks of a chain have the same period.
2.3 Special Properties
@@
In the paper, we consider both general $PD|T_{i}^{harm},jit=0,chains|objectives$ instances, as well as instances generated with a set of realistic properties. For example, we generate instances with specifically engineered chains and assignment of tasks to resources, so the instance models periodic communications over a network with a given topology. Some of these properties, which will be used in the following sections of the paper, are:
Limited number of resources:
@@
Harmonic processing times:
In this case, all values of processing times and periods are together in a harmonic set. This makes the single resource feasibility problem polynomial [7]. Graham notation for such case is $PD|(T_{i},p_{i})^{harm},jit=0,chains|objectives$.
Topology:
In this case, the underlying network topology is organized in such a way that the support graph of the instance has some special structure (e.g., a line or a tree).
Equal processing time on a chain:
In this case, inspired by the most common application of communication protocols, point-to-point communications have the same duration since the packets have the same length and communication links have the same speed. In such case $p_{i}^{k}=p^{k}$.
Acyclic:
In this setting, we assume the support graph contains no directed cycle. Proofs related to this case are described in Section 3.
Bottleneck resource:
A special case where there is at least one resource $b$ such that all chains go through $b$.
Scattering:
A special case of an acyclic instance with a bottleneck resource $b$ where all the first tasks of the chains are assigned to $b$ (as in the context of parallel computing [40]).
@@ 2.4 Structural Properties
In this subsection, we recall some structural properties useful to understand our algorithms and proofs of the next sections. First, we describe how to build a feasible schedule from a core schedule that only satisfies the resource constraints. Then, we describe the analogy introduced in [11] and formally established in [13], of a schedule on one resource and a 2D packing problem. The packing view is a crucial inspiration behind the warm-start approach described later.
2.4.1 Building a Schedule from Independent Resource Schedules
Let us consider a resource $m$ and the subset of tasks $\mathcal{J}_{m}$ assigned to $m$. We first define the following concept:
Definition 12 (Core schedule).
We call core schedule on resource $m$ any schedule $\sigma$ of the tasks $\mathcal{J}_{m}$ that satisfies the resource constraint, and such that
$\forall J_{i}^{k}\in\mathcal{J}_{m},\sigma_{i}^{k}<T^{k}$ (11)
Obviously, if $t$ is a schedule of tasks $\mathcal{J}_{m}$ that satisfies the resource constraints on $m$, then the following $\sigma$ is a core schedule on $m$.
$\sigma_{i}^{k}=t_{i}^{k}\mod T^{k},\quad\forall J_{i}^{k}\in\mathcal{J}_{m}$ (12)
Note that if $t$ is not integer, then the $\mod$ operation here is defined as follows: $t_{i}^{k}=q\cdot T^{k}+\sigma_{i}^{k}$, with an integer $q$, and $\sigma_{i}^{k}\in[0,T^{k})$.
We now prove that from a core schedule $\sigma$ on each resource, we can build a feasible schedule with core $\sigma$ that satisfies the resource and the precedence constraint and minimizes the degeneracy among all schedules with core $\sigma$. Indeed, we can satisfy the precedence constraints (3) of each chain $k$ by postponing a subsequent task in the chain to the next period whenever needed to satisfy the precedence constraint. For this purpose, we define Algorithm 1.
Input : Chains, tasks, resources, core schedule $\sigma$ on each resource
Output : Schedule $t_{i}^{k}=\sigma_{i}^{k}+q_{i}^{k}T^{k}$, with $q_{i}^{k}\in\mathbb{N}$ for each task $J_{i}^{k}$
for $k\leftarrow 1$ to $\mu$ do
       $q_{1}^{k}\leftarrow 0$
      for $i\leftarrow 1$ to $n^{k}-1$ do
             $q_{i+1}^{k}\leftarrow\left\{\begin{array}[]{lll}q_{i}^{k}&if&\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}\\ q_{i}^{k}+1&if&\sigma_{i+1}^{k}<\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}+T^{k}\\ q_{i}^{k}+2&&otherwise\end{array}\right.$
       end for
end for
Algorithm 1 Building a feasible schedule from a core schedule
The following lemma establishes that the schedule built by Algorithm 1 is indeed feasible with respect to the precedence constraints and has minimal degeneracy among schedules with core $\sigma$.
Lemma 1.
Assume that we are given for each $m\in\{1,\ldots,M\}$, a core schedule $\sigma$ of tasks in $\mathcal{J}_{m}$. Algorithm 1 outputs a schedule $t$ satisfying the precedence constraints and for which the degeneracy of each chain $\mathcal{C}^{k}$ is minimal among all schedules with core schedule $\sigma$.
Proof.
Notice first that any schedule $\hat{t}$ with core $\sigma$ satisfies, for each chain $\mathcal{C}^{k}$ and task $J_{i}^{k}$:
$\hat{t}_{i}^{k}=\sigma_{i}^{k}+\gamma_{i}^{k}T^{k}$ (13)
with $\gamma_{i}^{k}\in\mathbb{N}$, since $\hat{t}_{i}^{k}\mod T^{k}=\sigma_{i}^{k}$. So the schedule computed by Algorithm 1 where $t_{i}^{k}=\sigma_{i}^{k}+q_{i}^{k}T^{k}$ has indeed core $\sigma$.
Consider now the start-to-end latency of chain $\mathcal{C}^{k}$ for any schedule $\hat{t}$ with core $\sigma$:
$S^{k}(\hat{t})=\hat{t}_{n^{k}}^{k}+p_{n^{k}}^{k}-\hat{t}_{1}^{k}=\sigma_{n^{k}}^{k}+p_{n^{k}}^{k}-\sigma_{1}^{k}+(\gamma_{n^{k}}^{k}-\gamma_{1}^{k})T^{k}$ (14)
so:
$D^{k}(\hat{t})=\gamma_{n^{k}}^{k}-\gamma_{1}^{k}+\left\lceil\frac{\sigma_{n^{k}}^{k}+p_{n^{k}}^{k}-\sigma_{1}^{k}}{T^{k}}\right\rceil-1$ (15)
So, among schedules with core $\sigma$, the degeneracy just depends on the difference $\gamma_{n^{k}}^{k}-\gamma_{1}^{k}$.
We now prove that the schedule computed by Algorithm 1 satisfies the precedence constraint and that $q_{n^{k}}^{k}-q_{1}^{k}$ is minimal for any chain $\mathcal{C}^{k}$.
Consider a task $J_{i}^{k}$ and its subsequent task $J_{i+1}^{k}$. Notice that the corresponding precedence constraint on any schedule with core $\sigma$ is expressed as follows:
$\hat{t}_{i+1}^{k}\geq\hat{t}_{i}^{k}+p_{i}^{k}\quad\Leftrightarrow\quad\gamma_{i+1}^{k}-\gamma_{i}^{k}\geq\left\lceil\frac{\sigma_{i}^{k}+p_{i}^{k}-\sigma_{i+1}^{k}}{T^{k}}\right\rceil$ (16)
Notice that Algorithm 1 distinguishes three cases. If $\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}$ then:
$q_{i+1}^{k}-q_{i}^{k}=0=\left\lceil\frac{\sigma_{i}^{k}+p_{i}^{k}-\sigma_{i+1}^{k}}{T^{k}}\right\rceil,$ (17)
so the schedule $t$ satisfies equation (16).
Similarly, if $\sigma_{i+1}^{k}<\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}+T^{k}$, then:
$q_{i+1}^{k}-q_{i}^{k}=1=\left\lceil\frac{\sigma_{i}^{k}+p_{i}^{k}-\sigma_{i+1}^{k}}{T^{k}}\right\rceil,$ (18)
and equation (16) holds.
In the latter case, we necessarily have $\sigma_{i}^{k}+p_{i}^{k}<\sigma_{i+1}^{k}+2T^{k}$, since $\sigma_{i}^{k}$ and $p_{i}^{k}$ are lower than $T^{k}$. Then:
$q_{i+1}^{k}-q_{i}^{k}=2=\left\lceil\frac{\sigma_{i}^{k}+p_{i}^{k}-\sigma_{i+1}^{k}}{T^{k}}\right\rceil,$ (19)
and equation (16) holds.
Moreover in the three subcases, for any other schedule $\hat{t}$ with core $\sigma$ we have
$\gamma_{i+1}^{k}-\gamma_{i}^{k}\geq q_{i+1}^{k}-q_{i}^{k}$ (20)
Hence, by summing these inequalities:
$\gamma_{n^{k}}^{k}-\gamma_{1}^{k}\geq q_{n^{k}}^{k}-q_{1}^{k}.$ (21)
So, the degeneracy of each chain $\mathcal{C}^{k}$ is minimal for the schedule build by Algorithm 1.
(a) The network consists of four network nodes A, B, C, D in line topology connected with full-duplex links. Ten messages are sent from the leftmost network node A to the rightmost node D via three blue links (i.e., resources). Corresponding chains $\mathcal{C}^{1}-\mathcal{C}^{10}$, each consisting of three tasks $J_{i}^{k}$, are to be scheduled on resources 1, 2, 3 (assignment of tasks to resources $\pi_{i}^{k}=i$ is shown by dotted ellipses). For simplicity, no messages are sent via resources 4, 5, 6 in the opposite direction.
(b) Schedule of ten chains on three resources spanning over one hyperperiod. Tasks of the same chain have the same color. Total degeneracy is 2. Start-to-end latency $S^{k}$ is shown for light blue $\mathcal{C}^{1}$ and yellow $\mathcal{C}^{3}$ chains, which both have degeneracy $D^{k}=1$.
Figure 2: Illustrative communication case and associated schedules for instance with periods $\tau(1)=40,\tau(2)=80,\tau(3)=240$. Ten chains are to be scheduled; $\mathcal{C}^{1}$ and $\mathcal{C}^{2}$ (dark and light blue) have period 40 (thus $T^{1}=T^{2}=\tau(1)$, meaning $\beta^{1}=\beta^{2}=1$), $\mathcal{C}^{3}-\mathcal{C}^{7}$ (e.g., yellow) have period 80, and $\mathcal{C}^{8}-\mathcal{C}^{10}$ (e.g., cyan) have period 240. Notice that all the messages are sent from the leftmost network node to the rightmost one; communication links in the other direction are not used. Utilization $U_{m}=0.96$ holds for all resources.
Such schedules, shown within a single hyperperiod, can be illustrated as in Fig. 2(b). There, 10 chains are delivered using a sequence of resources 1, 2, and 3, as is shown in Fig. 2(a). Precedence constraints for chains hold, and no overlaps occur on either resource.
@@ 2.4.2 Single Resource Core Schedule Coordinates
Consider a single resource $m$ and a feasible core schedule $\sigma$ of the independent periodic tasks of $\mathcal{J}_{m}$. Let $o_{m}$ be the start time of the first task $J_{f}^{k}$ on resource $m$ with the least period $w$. We consider $o_{m}$ as the offset for resource $m$ from 0. So we define the shifted core schedule $s_{i}^{k}$ for each task $J_{i}^{k}\in\mathcal{J}_{m}$ such that:
@@
We call time interval $[o_{m},o_{m}+\hat{H})$ the “observation interval”. Notice that for any integer $v\geq 1$, the schedule $\sigma$ of time interval $[o_{m}+(v-1)\hat{H},o_{m}+v\hat{H})$ on resource $m$ is the same as the schedule of the observation interval shifted by $(v-1)\hat{H}$ time units. This offset may differ for different resources.
@@ 2.4.3 Associated HD2D Packing Problem
To each task $J_{i}^{k}\in\mathcal{J}_{m}$ we associate a rectangle $R_{i}^{k}$ of width $\ell_{i}^{k}=p_{i}^{k}$ and height $0pt{i}{k}=\frac{\hat{H}}{T^{k}}$, to fit in a bin of width $w$ and height $H=\frac{\hat{H}}{w}$ without overlapping rectangles. Moreover, the packing has to be height divisible, which means that the vertical coordinate of each rectangle is an integer multiple of its height.
Theorem 1 ([13]).
There exists a feasible schedule of tasks of $\mathcal{J}_{m}$ on machine $m$ if and only if there exists a height divisible packing of the rectangles of the associated instance in a bin of width $w$ and height $H=\frac{\hat{H}}{w}$.
(a) Schedule of a single observation interval of resource 2 arranged in rows of length 40.
(b) Height-divisible 2D packing of resource 2.
Figure 3: Observation interval and packing perspective of a schedule of resource 2 from Fig. 2(b). The offset of the shifted core schedule is $o_{2}=40$.
Paper [13] gives the formal correspondence between the start time of a task in the core schedule $s_{i}^{k}$ and coordinates of the bottom left corner of the associated rectangle. To give an intuition of this correspondence, we develop it on an example in Fig. 3. The main idea is to arrange the periodic schedule in rows and shuffle them to produce the packing.
In Fig. 3(a), the single “observation interval” associated with the schedule of resource 2 from the motivational example of Fig. 2(b) is shown. Ten chains $\mathcal{C}^{1},\dots,\mathcal{C}^{10}$ pass through the resource. $\mathcal{C}^{1},\mathcal{C}^{2}$ have the shortest period $\tau(1)=40$. Chains $\mathcal{C}^{3}-\mathcal{C}^{7}$ have a period equal to 80; the rest have a period equal to 240, which is the hyperperiod. Origin 00 is fixed by the start time of chain $\mathcal{C}^{1}$ (resource’s offset is $o_{2}=40$), and we stack each interval of length $w=40$ on top of the previous one. So row $v$ describes the schedule of interval $[v\cdot w,(v+1)w)$. Horizontal coordinate $u$ corresponds to the relative position of the task’s occurrence in the interval.
@@
Figure 3(b) shows the corresponding height-divisible packing view of the same solution. Rectangles are described with coordinates of their bottom-left corner $(x,y)$. The packing is obtained by interchanging the rows as follows: we swap the two-digit row numbers in Figure 3(a) such that the most significant digit becomes the least significant digit and vice versa. Finally, we obtain Figure 3(b) by ordering the rows according to their new two-digit numbers. For example, the row with $\mathcal{C}^{6}$ has number $10_{3}1_{2}=$ 1 in Figure 3(a) and after the swap, the same row has number $31_{2}0_{3}=$ 3 in Figure 3(b). We can see the occurrences of the original tasks have been merged, and the dimensions of the resulting rectangles correspond to the rectangles of the associated packing instance. For example, tasks 1-2 with the shortest period are transformed into rectangles spanning the entire height of the bin. Tasks 3-7 merge three of their occurrences into a rectangle with height 3, and finally, rectangles associated with the longest-period tasks 8-10 have a height equal to 1.
@@ 3 Complexity Analysis
In this section, we consider related complexity results to explain the complexity of our problems. First, we know that scheduling periodic tasks with harmonic periods on a single resource is a strongly NP-hard problem [6] whereas it becomes polynomial if processing times and periods belong to the harmonic set [7]. So, minimizing one of our objectives seems at least as difficult as this problem and is still NP-hard unless very strong assumptions are made. In the rest of the section, we show some special polynomial cases of our problem and present a strongly NP-hard subproblem.
3.1 Two Resources with Chains of Length Two.
In this subsection, we assume that there are only two resources and that chains either start with a task on resource 1 and end with a task on resource 2, or vice versa.
3.1.1 Upper Bounds for the Two Resource Case
We prove that if we build a schedule as in Algorithm 1 from a core schedule on each resource, then the degeneracy is bounded.
Property 1.
Assume that $\mathcal{C}^{k}$ is a chain of length two on two different resources. A schedule $t$ built according to Algorithm 1 satisfies: if $p_{1}^{k}+p_{2}^{k}\leq T^{k}$ then $D^{k}\leq 1$ else $D^{k}\leq 2$.
Proof.
We consider a chain $\mathcal{C}^{k}$, and we assume without loss of generality that it starts on the first resource. Let us consider $\sigma$ a core schedule satisfying the resource constraints, and assume that $\sigma_{i}^{k}<T^{k}$ for $i\in\{1,2\}$. We can observe that due to the fact that the processing time of each task is not greater than $T^{k}$, we have the following inequalities:
$0\leq p_{1}^{k}+p_{2}^{k}<2T^{k}$ (22)
Notice that in any case of the construction given in Algorithm 1, $t_{1}^{k}=\sigma_{1}^{k}$. Let us consider the different cases of computation of $t_{2}^{k}$ applied in Algorithm 1. For each case, we show the bounds of degeneracy.
Case 1: $\sigma_{1}^{k}+p_{1}^{k}\leq\sigma_{2}^{k}$. In this case Algorithm 1 sets $t_{2}^{k}=\sigma_{2}^{k}$.
$S^{k}=\sigma_{2}^{k}+p_{2}^{k}-\sigma_{1}^{k}<\sigma_{2}^{k}+p_{2}^{k}$ (23)
$\Rightarrow S^{k}<2T^{k}\quad\Rightarrow\quad D^{k}\leq 1$ (24)
Case 2: $\sigma_{2}^{k}<\sigma_{1}^{k}+p_{1}^{k}\leq\sigma_{2}^{k}+T^{k}$. Then Algorithm 1 sets $t_{2}^{k}=\sigma_{2}^{k}+T^{k}$. By Eq. 22:
$S^{k}=\sigma_{2}^{k}+p_{2}^{k}+T^{k}-\sigma_{1}^{k}<p_{1}^{k}+p_{2}^{k}$ (25)
$\Rightarrow S^{k}<2T^{k}\quad\Rightarrow\quad D^{k}\leq 1$ (26)
Notice that in the special case where $p_{1}^{k}+p_{2}^{k}\leq T^{k}$, then $D^{k}=0$.
Case 3: $\sigma_{2}^{k}+T^{k}<\sigma_{1}^{k}+p_{1}^{k}$. Then Algorithm 1 sets $t_{2}^{k}=\sigma_{2}^{k}+2T^{k}$. In this case:
$\sigma_{2}^{k}+p_{2}^{k}-\sigma_{1}^{k}<p_{1}^{k}+p_{2}^{k}-T^{k}$ (27)
$\Rightarrow S^{k}=\sigma_{2}^{k}+p_{2}^{k}+2T^{k}-\sigma_{1}^{k}<p_{1}^{k}+p_{2}^{k}+T^{k}$ (28)
Now in any case from Eq. 22,
$S^{k}<3T^{k}\quad\Rightarrow\quad D^{k}\leq 2$ (29)
Notice that in the special case where $p_{1}^{k}+p_{2}^{k}\leq T^{k}$, then we would have $S^{k}\leq 2T^{k}$, so that $D^{k}\leq 1$.
Corollary 1.
In the case of processing times and periods belonging to a harmonic set, building a schedule on two resources of chains of length two with $D_{\max}\leq 1$ can be done in polynomial time.
Proof.
Notice that if processing times are harmonic, then the processing time of a task is at most half of the period. Hence, we can easily establish that in this case, $p_{1}^{k}+p_{2}^{k}\leq T^{k}$. This implies that building two independent periodic schedules on each resource in polynomial time with the algorithm of [7] and then applying Algorithm 1 leads to a schedule in which for each chain $\mathcal{C}^{k}$, $D^{k}\leq 1$ according to Property 1. ∎
3.1.2 Scattering on a Line of Two Resources with a Single Period